OpenStudy (anonymous):

What is the relationship between the sine and cosine of complementary angles?

OpenStudy (johnweldon1993):

\[\large sin(\theta )= cos(90^\circ - \theta)\] And \[\large cos(\theta )= sin(90^\circ - \theta)\] It works because sine and cosine are "co-functions"